OpenStudy (anonymous):

is the inverse function of ^3 square root 2x-5 equal to f^-1 = 2x^3+5 ??

OpenStudy (mr.math):

Write \(x=(2y-5)^{1/3}\) and solve for y, \[x^3=2y-5 \implies 2y=x^3+5 \implies y=\frac{1}{2}(x^3+5)=f^{-1}(x).\]